# Enhancing the Performance of Aspect-Based Sentiment Analysis Systems

### v2

@@ Abstract
Aspect-based sentiment analysis aims to predict sentiment polarity with fine granularity. While Graph Convolutional Networks (GCNs) are widely utilized for sentimental feature extraction, their naive application for syntactic feature extraction can compromise information preservation. This study introduces an innovative edge-enhanced GCN, named SentiSys, to navigate the syntactic graph while preserving intact feature information, leading to enhanced performance. Specifically,we first integrate a bidirectional long short-term memory (Bi-LSTM) network and a self-attention-based transformer. This combination facilitates effective text encoding, preventing the loss of information and predicting long dependency text. A bidirectional GCN (Bi-GCN) with message passing is then employed to encode relationships between entities. Additionally, unnecessary information is filtered out using an aspect-specific masking technique. To validate the effectiveness of our proposed model, we conduct extensive evaluation experiments and ablation studies on four benchmark datasets. The results consistently demonstrate improved performance in aspect-based sentiment analysis when employing SentiSys. This approach successfully addresses the challenges associated with syntactic feature extraction, highlighting its potential for advancing sentiment analysis methodologies.
@@
Performance Evaluation, Aspect-based Sentiment Analysis Systems, Graph Convolutional Networks
@@ 1 Introduction
Aspect-based sentiment analysis [1] aims to classify sentences in aspect granularity. The sentiments of each aspect can be different even in the same sentence. Aspect-level sentiment analysis has been widely used in real-world applications. Suppose there is such a review comment among the restaurant reviews: “Although the menu is limited, the friendly staff provided us with a nice night.” If we pay attention to the aspect of “menu”, the sentiment is negative (“limited”); when it comes to “staff”, we can extract the opposite sentiment, i.e., positive (“friendly”). Such multi-aspect can introduce multi-sentiments in the same sentence, resulting in difficulty of sentiment prediction. To solve such difficulty, aspect-based sentiment analysis is attracting attention to improve services [2].
Figure 1: An example of spurious associations of a customer review system.
@@
Neural networks are widely used for aspect-based sentiment analysis because of their promising results, especially recurrent neural networks (RNNs) [3] and the Transformer [4] have demonstrated their sequence modeling capabilities. [5, 6, 7, 8, 9]. However, these models are not affordable to establish precise dependencies between word pairs and aspects. Suppose such a sentence: “The dish looks mediocre but tastes surprisingly wonderful!” Although the aspect “dish” is linked to two opposite sentiments, “looks mediocre” and “tastes surprisingly wonderful”, we can easily understand that the latter is more tightly linked to the “dish”. However, this judgment may be difficult for these models due to the closer distance of the former to the “dish”. Convolutional neural networks (CNNs) [10] are studied to make predictions based on phrases that express sentiments more accurately. However, phrases generally have longer dependencies on the subject. For example, “tastes surprisingly wonderful” may express a positive sentiment more accurately than just “wonderful”. CNNs are not affordable for such long dependencies because of the local receptive field limitation [11].
Graph convolutional networks (GCNs) [12, 13] can decompose sentences into a tree structure composed of entities. Such a tree structure constructs a more direct path between aspects and related words. Therefore, GCNs resolve the long-distance dependency problem and facilitate the learning of the representation of nodes and their positional relationships [14]. However, the general GCNs lack the edge information. GCNs cannot capture the syntax and semantic relationships due to the lack of edge information resulting in poor performance [15, 16]. Figure 1 shows examples that illustrate these complex relationships in customer reviews. The aspect “service” is associated with “never had”, and its sentiment is positive in past reviews. However, in a new customer review, its sentiment is not positive. The network cannot determine the sentiment by relying on the phrase “never had”.
@@
This study is an extension of our previous work [17], introducing an innovative edge-enhanced graph convolutional network (GCN) named SentiSys, designed to elevate the effectiveness of aspect-based sentiment analysis systems. In our approach, we leverage bidirectional long short-term memory (Bi-LSTM) to extract node word embeddings, and Bi-GCN is employed to model word dependencies within a tree structure, thereby enhancing word representations. The integration of these components facilitates accurate association of sentiment words with the target aspect, resulting in a notable performance improvement. The robustness of the proposed SentiSys model is substantiated through extensive evaluation experiments, confirming its high-performance capabilities. The key contributions of this work are encapsulated in three primary aspects, providing a comprehensive enhancement to aspect-based sentiment analysis methodologies.
@@
  • Architecture based on neural networks: The SentiSys incorporates a powerful combination of Bi-LSTM, a transformer, and an edge-enhanced Bi-GCN. This intricate setup leverages linguistic syntax to construct a graph, which, in turn, serves as the foundation for extracting and representing textual content through a sophisticated GCN. This approach underscores the importance of effectively harnessing these cutting-edge technologies to decode sentiment and meaning within the text, making it a notable advancement in sentiment analysis.
@@
  • Effective semantic edges: Employing a dependency tree not only reduces the gap between aspects and their related words but also establishes a syntactic pathway within the text. Additionally, SentiSys takes into account the edge information within syntactic dependency trees. This valuable data enhances the syntactic connection between associated words and the specific target aspect.
@@ 2.1 Aspect-based Sentiment Analysis Systems
Conventional sentiment analysis methodologies typically involve training classifiers utilizing features such as bag-of-words and sentiment dictionaries [18]. However, both rule-based [19] and statistical-based strategies [20] require handcrafted features, revealing the limitations of these naive sentiment analysis methods, particularly when confronted with the complex intricacies inherent in customer review analysis. These shortcomings include challenges in capturing nuanced sentiments related to specific aspects of interest.
Motivated by the inadequacies of traditional sentiment analysis, the research landscape has progressively shifted towards aspect-based sentiment analysis. Aspect-based sentiment analysis encompasses four key tasks [21]: aspect term sentiment analysis (ATSA) [22], aspect category sentiment analysis (ACSA) [23], aspect term extraction (ATE) [24], and aspect category extraction (ACE) [25]. ATSA involves training classifiers to predict the sentiment polarity associated with specified aspect terms, addressing the need for a more granular understanding of sentiment expressions within the context of specific elements mentioned in reviews [26]. ACSA, on the other hand, aims to predict sentiment polarity concerning predefined sentiment categories, providing insights into broader sentiment trends related to predetermined aspects [27]. ATE and ACE, conceptualized as entity extraction challenges, strive to solve the intricate problem of identifying and understanding aspects within reviews. ATE focuses on extracting precise aspect terms, such as the previously mentioned example of “dish,” aiming to capture specific entities of interest [28]. Conversely, ACE aims to identify and categorize broader aspects, like the “food” category, encompassing terms like “dish.”
In addressing these tasks, the motivation lies in achieving a more nuanced and targeted comprehension of sentiment expressions within diverse contexts. By refining the understanding of sentiments associated with specific aspects, aspect-based sentiment analysis contributes to enhanced accuracy and applicability in sentiment analysis practices [29]. However, it is essential to acknowledge potential limitations, such as the need for robust methods to handle ambiguous language and context, which opens avenues for further research in advancing the effectiveness of aspect-based sentiment analysis methodologies.
@@ 2.2 RNN-based Sentiment Analysis Systems
The recent remarkable advances in neural networks have attracted the attention of researchers for applications [30, 31, 32, 33, 34]. The models proposed in [35] for target-dependent sentiment classification leverage neural networks. Specifically, the LSTM architecture, originally designed for sequence modeling, proves effective for aspect-based sentiment analysis tasks in an autoregressive manner. Building upon the LSTM framework, this approach enhances the architecture by incorporating target-specific information. Referred to as target-dependent LSTM (TD-LSTM), this model skillfully captures the dynamic interplay between a target word and its contextual words, discerningly selecting pertinent segments of the context to deduce sentiment polarity towards the target. The model is amenable to end-to-end training through standard backpropagation, employing the cross-entropy error of supervised sentiment classification as its loss function. While the TD-LSTM approach enhances sentiment classification by considering target-specific information, the nuanced understanding of the interplay between a target word and its contextual words remains a complex task. This limitation may impact the model’s ability to capture subtle sentiment nuances and context-specific variations accurately.
@@
Of significance in this domain is the Transformer [4], an innovative architecture explicitly designed to excel in capturing intricate long-range dependencies within textual data. The Transformer’s demonstrated effectiveness positions it as a compelling choice for tasks requiring a robust treatment of extensive contextual relationships.
@@ 2.3 GCN-based Sentiment Analysis Systems
Unlike RNN-based sentiment analysis systems, GCN-based systems seek to represent language data as graphs, where words or phrases are nodes and relationships between them are edges. This representation allows for a more nuanced understanding of contextual dependencies, capturing intricate connections that may be challenging for traditional RNNs [39].
@@
Several notable advancements in aspect-based sentiment analysis using GCNs have been introduced. ASGCN [40] showcases remarkable proficiency in capturing sentiment polarity across various aspects in a text. This is achieved through its adept modeling of relationships between aspects and their respective words utilizing GCNs. However, the efficacy of ASGCN is contingent on precise dependency tree parsing. Instances of inaccurate parsing pose a potential challenge, introducing a risk of suboptimal modeling. This directly undermines the fidelity of relationships between aspects and words. Consequently, the overall performance of ASGCN is adversely affected, especially within the domain of aspect-based sentiment analysis. AEGCN [41] seamlessly integrates GCNs with attention mechanisms, specifically embracing multi-head interactive attention and multi-head self-attention. The prowess of AEGCN crucially hinges on the precise encoding of attention. Any deviations or inaccuracies in this intricate process pose a risk of suboptimal modeling, directly impacting the relationships within the graph. SK-GCN [42] combines a syntactic dependency tree and GCN to model the syntax relationship more precisely. However, SK-GCN encounters a significant limitation attributed to the graph structure of commonsense knowledge. Specifically, the model faces challenges arising from an insufficient graph structure when representing commonsense knowledge. This limitation can impact the model’s ability to effectively leverage commonsense knowledge for aspect-level sentiment analysis. AGCN [43] aggregates the feature information of neighboring nodes to make the node representations related to each other. However, the effectiveness of AGCN heavily depends on the accuracy and completeness of the graph representation. In scenarios where the graph structure does not accurately capture the relationships within the data, AGCN may struggle to generate meaningful and representative features, leading to suboptimal performance.
@@
Unlike previous studies, we present a novel integration ofBi-LSTM, Transformer, and GCNs to advance aspect-based sentiment analysis. SentiSys strategically leverages the strengths of each component: Bi-LSTM serves as a proficient feature extractor, capturing intricate patterns and semantic nuances within the text. Meanwhile, the Transformer excels at capturing long-distance dependencies, allowing our model to grasp the contextual relationships crucial for comprehensive sentiment analysis. Introducing GCN with a dependency parsing layer further enriches the model’s capabilities by incorporating syntactic information, enabling a more fine-grained understanding of the syntax relations within the text. By combining these components, SentiSys aims to provide a holistic and nuanced perspective on aspect-based sentiment analysis. This innovative amalgamation addresses the limitations of prior studies, fostering a more comprehensive and accurate sentiment analysis methodology.
3 SentiSys Model Description
In this section, we mainly describe the proposed SentiSys in seven aspects. The word embedding layer creates low-dimensional representations for each token in a vocabulary; an Bi-LSTM aims to extract features of sentences; a Transformer model is used to extract contextual features from a global perspective; a dependency parsing layer creates a tree structure of context grammar; an Bi-GCN processes the generated tree structure; an aspect-specific masking layer is used to reduce redundancy and enhance the accuracy; the sentiment classification layer classifies the sentiment of a given sentence. Figure 2 shows an overview architecture of the proposed SentiSys for the aspect-based sentiment analysis.
@@ 3 SentiSys Model Description
Figure 2: The architectural overview of the proposed SentiSys for aspect-based sentiment analysis comprises the following key components: (a) The initial word embedding layer parses sentences and embeds words into a vector space, forming the foundation for further analysis. (b) The Bi-LSTM network enables the capture of contextual information in a sentence for a deeper understanding of word relationships. (c) The transformer layer with a self-attention mechanism can explore the interrelationships and global word features in lengthy text, enhancing contextual understanding. (d) The dependency parsing layer can recognize the dependencies among the words in a sentence and construct a dependency tree to help explain the grammatical structure. (e) The Bi-GCN propagates information among the nodes of the parsed dependency tree using message passing, thus effectively modeling word relationships. (f) The aspect-specific masking technique refines the hidden representation by separating the information related to the aspect of interest, which reduces redundancy and improves the accuracy of aspect-based sentiment analysis. (g) The final sentiment classification layer is dedicated to classifying the sentiment of a given sentence into predefined categories, typically categorizing it as positive, negative, or neutral.
@@ 3.1 Word Embedding Layer
To tokenize a sentence consisting of $n$ words with an $m$-word aspect, we represent it as a sequence denoted as ${\bf S}$ in the following manner:
${\bf S}=\left\{{\bf w}_{1},{\bf w}_{2},\cdot\cdot\cdot,{\bf w}_{\gamma},\cdot\cdot\cdot,{\bf w}_{\gamma+m},\cdot\cdot\cdot,{\bf w}_{n}\right\},$ (1)
where ${\bf w}_{t}\in{\bf S}$ and $\gamma$ indicate a word token and starting index of an aspect, respectively. Each word token is then transformed into a low-dimensional vector denoted as ${\bf e}_{t}\in\mathbb{R}^{d_{w}}$, where $d_{w}$ denotes the dimension of the word embedding.
@@ 3.2 Bi-LSTM Network
Word embedding embeds the words in the sentence into low-dimensional vectors, and these embedding vectors are used as inputs to the Bi-LSTM network [44]. The utilization of Bi-LSTM allows for the extraction of contextual information from sentences. As mentioned in [45], the Bi-LSTM network comprises two components: the forward LSTM and the backward LSTM. In the forward LSTM, the hidden vector $\overrightarrow{\bf h}^{LSTM}_{t}$ at the current time step is computed by combining the input embedding ${\bf e}_{t}$ with $\overrightarrow{\bf h}^{LSTM}_{t-1}$ from the previous time step. The backward LSTM is constructed by utilizing ${\bf e}_{t}$ and $\overleftarrow{\bf h}^{LSTM}_{t+1}$ from the subsequent time step. Finally, the hidden vectors $\overrightarrow{\bf h}^{LSTM}_{t}$ and $\overleftarrow{\bf h}^{LSTM}_{t}$ are concatenated to create the overall hidden vector ${\bf h}^{LSTM}_{t}$. This process is illustrated as follows:

### v3

@@ Abstract
Aspect-based sentiment analysis predicts sentiment polarity with fine granularity. While graph convolutional networks (GCNs) are widely utilized for sentimental feature extraction, their naive application for syntactic feature extraction can compromise information preservation. This study introduces an innovative edge-enhanced GCN, named SentiSys, to navigate the syntactic graph while preserving intact feature information, leading to enhanced performance. Specifically,we first integrate a bidirectional long short-term memory (Bi-LSTM) network and a self-attention-based transformer. This combination facilitates effective text encoding, preventing the loss of information and predicting long dependency text. A bidirectional GCN (Bi-GCN) with message passing is then employed to encode relationships between entities. Additionally, unnecessary information is filtered out using an aspect-specific masking technique. To validate the effectiveness of our proposed model, we conduct extensive evaluation experiments on four benchmark datasets. The experimental results demonstrate enhanced performance in aspect-based sentiment analysis with the use of SentiSys.
@@
Aspect-based sentiment analysis systems Graph convolutional networks
@@ 1 Introduction
Aspect-based sentiment analysis [15] aims to classify sentences in aspect granularity. The sentiments of each aspect can be different even in the same sentence. Aspect-level sentiment analysis has been widely used in real-world applications. Suppose there is such a review comment among the restaurant reviews: “Although the menu is limited, the friendly staff provided us with a nice night.” If we pay attention to the aspect of “menu”, the sentiment is negative (“limited”); when it comes to “staff”, we can extract the opposite sentiment, i.e., positive (“friendly”). Such multi-aspect can introduce multi-sentiments in the same sentence, resulting in difficulty of sentiment prediction. To solve such difficulty, aspect-based sentiment analysis is attracting attention [2].
@@
Neural networks are widely used for aspect-based sentiment analysis because of their promising results, especially recurrent neural networks (RNNs) [38] and the Transformer [18, 30] have demonstrated their sequence modeling capabilities. However, these models are not affordable to establish precise dependencies between word pairs and aspects. Suppose such a sentence: “The dish looks mediocre but tastes surprisingly wonderful!” Although the aspect “dish” is linked to two opposite sentiments, “looks mediocre” and “tastes surprisingly wonderful”, we can easily understand that the latter is more tightly linked to the “dish”. However, this judgment may be difficult for these models due to the closer distance of the former to the “dish”. Convolutional neural networks (CNNs) [28, 36] are studied to make predictions based on phrases that express sentiments more accurately. However, phrases generally have longer dependencies on the subject. For example, “tastes surprisingly wonderful” may express a positive sentiment more accurately than just “wonderful”. CNNs are not affordable for such long dependencies because of the local receptive field limitation [14].
@@
Graph convolutional networks (GCNs) [34, 40] can decompose sentences into a tree structure composed of entities. Such a tree structure constructs a more direct path between aspects and related words. Therefore, GCNs resolve the long-distance dependency problem and facilitate the learning of the representation of nodes and their positional relationships [22]. However, the general GCNs lack the edge information. GCNs cannot capture the syntax and semantic relationships due to the lack of edge information resulting in poor performance [4]. This study introduces an innovative edge-enhanced graph convolutional network (GCN) named SentiSys, designed to elevate the effectiveness of aspect-based sentiment analysis systems. In our approach, we leverage bidirectional long short-term memory (Bi-LSTM) to extract node word embeddings, and Bi-GCN is employed to model word dependencies within a tree structure, thereby enhancing word representations. The integration of these components facilitates accurate association of sentiment words with the target aspect, resulting in a notable performance improvement. The robustness of the proposed SentiSys is substantiated through extensive evaluation, confirming its high-performance capabilities. The main contributions are as follows:
@@
  • Architecture based on neural networks: SentiSys integrates Bi-LSTM, a transformer, and an edge-enhanced Bi-GCN, leveraging linguistic syntax to construct a graph. This graph serves as the foundation for extracting and representing textual content through a sophisticated GCN. This approach highlights the significance of harnessing these advanced technologies to decode sentiment and meaning within text, marking a notable advancement in sentiment analysis.
@@
  • Effective semantic edges: Using a dependency tree minimizes the gap between aspects and related words while establishing a syntactic pathway within the text. SentiSys also considers edge information within syntactic dependency trees, enhancing syntactic connections between associated words and the target aspect.
@@ 2.1 Aspect-based Sentiment Analysis Systems
Conventional sentiment analysis methodologies typically involve training classifiers utilizing features such as bag-of-words and sentiment dictionaries. However, both rule-based [8] and statistical-based strategies [13] require handcrafted features, revealing the limitations of these naive sentiment analysis methods, particularly when confronted with the complex intricacies inherent in customer review analysis. These challenges arise in capturing nuanced sentiments regarding specific aspects.
Motivated by the inadequacies of traditional sentiment analysis, the research landscape has progressively shifted towards aspect-based sentiment analysis. Aspect-based sentiment analysis encompasses four key tasks [3]: aspect term sentiment analysis (ATSA), aspect category sentiment analysis (ACSA), aspect term extraction (ATE), and aspect category extraction (ACE). ATSA involves training classifiers to predict the sentiment polarity associated with specified aspect terms, addressing the need for a more granular understanding of sentiment expressions within the context of specific elements mentioned in reviews [24]. ACSA, on the other hand, aims to predict sentiment polarity concerning predefined sentiment categories, providing insights into broader sentiment trends related to predetermined aspects [20]. ATE and ACE, conceptualized as entity extraction challenges, strive to solve the intricate problem of identifying and understanding aspects within reviews. ATE focuses on extracting precise aspect terms, such as the previously mentioned example of “dish,” aiming to capture specific entities of interest [1]. Conversely, ACE aims to identify and categorize broader aspects, like the “food” category, encompassing terms like “dish.”
In addressing these tasks, the motivation lies in achieving a more nuanced and targeted comprehension of sentiment expressions within diverse contexts. By refining the understanding of sentiments associated with specific aspects, aspect-based sentiment analysis contributes to enhanced accuracy and applicability in sentiment analysis practices [9]. Recognizing limitations is crucial, particularly in developing robust methods to address ambiguous context, thereby prompting further research to enhance the efficacy of aspect-based sentiment analysis.
@@ 2.2 RNN-based Sentiment Analysis Systems
The recent remarkable advances in neural networks have attracted the attention of researchers for applications [16, 17, 37]. The models proposed in [29] for target-dependent sentiment classification leverage neural networks. Specifically, the LSTM architecture, originally designed for sequence modeling, proves effective for aspect-based sentiment analysis tasks in an autoregressive manner. Building upon the LSTM framework, this approach enhances the architecture by incorporating target-specific information. Referred to as target-dependent LSTM (TD-LSTM), this model skillfully captures the dynamic interplay between a target word and its contextual words, discerningly selecting pertinent segments of the context to deduce sentiment polarity towards the target. The model is amenable to end-to-end training through standard backpropagation, employing the cross-entropy error of supervised sentiment classification as its loss function. While the TD-LSTM approach enhances sentiment classification by considering target-specific information, the nuanced understanding of the interplay between a target word and its contextual words remains a complex task. This limitation affects model accuracy in capturing subtle sentiment nuances and context-specific variations.
@@ 2.3 GCN-based Sentiment Analysis Systems
Unlike RNN-based sentiment analysis systems, GCN-based systems seek to represent language data as graphs, where words or phrases are nodes and relationships between them are edges. This representation enables a nuanced understanding of contextual dependencies, capturing intricate connections challenging for traditional RNNs [31].
@@
Several notable advancements in aspect-based sentiment analysis using GCNs have been introduced. ASGCN [33] showcases remarkable proficiency in capturing sentiment polarity across various aspects in a text. This is achieved through its adept modeling of relationships between aspects and their respective words utilizing GCNs. However, the efficacy of ASGCN is contingent on precise dependency tree parsing. Instances of inaccurate parsing pose a potential challenge, introducing a risk of suboptimal modeling. This directly undermines the fidelity of relationships between aspects and words. Consequently, the overall performance of ASGCN is adversely affected, especially within the domain of aspect-based sentiment analysis.
@@
Unlike prior research, we introduce a novel fusion of Bi-LSTM, Transformer, and GCNs to advance aspect-based sentiment analysis. SentiSys strategically utilizes each component’s strengths: Bi-LSTM acts as an adept feature extractor, capturing intricate patterns and semantic nuances. Meanwhile, the Transformer excels at capturing long-range dependencies, enabling our model to understand crucial contextual relationships. The introduction of GCN with a dependency parsing layer enriches the model by incorporating syntactic information, facilitating a more nuanced understanding of text syntax. By combining these elements, SentiSys aims to offer a comprehensive perspective on aspect-based sentiment analysis, addressing previous limitations and fostering a more accurate sentiment analysis methodology.
3 SentiSys
@@
Figure 1: SentiSys for aspect-based sentiment analysis includes the following components: (a) Initial word embedding layer for parsing and embedding words into vectors. (b) Bi-LSTM network captures contextual information for deeper word relationship understanding. (c) Transformer layer with self-attention mechanism explores global word features, enhancing contextual understanding. (d) Dependency parsing layer recognizes word dependencies, constructing a dependency tree to explain grammatical structure. (e) Bi-GCN propagates information among tree nodes, effectively modeling word relationships. (f) Aspect-specific masking refines hidden representation, reducing redundancy and improving accuracy. (g) Sentiment classification layer categorizes sentence sentiment into predefined categories, such as positive, negative, or neutral.
SentiSys consists of a word embedding layer for token representation, a Bi-LSTM for sentence feature extraction, a Transformer for global contextual feature extraction, a dependency parsing layer for context grammar tree creation, a Bi-GCN for processing the tree structure, an aspect-specific masking layer to reduce redundancy and enhance accuracy, and a sentiment classification layer for sentence sentiment classification. Figure 1 illustrates the architecture of SentiSys for aspect-based sentiment analysis.
@@ 3.1 Word Embedding Layer
To tokenize a sentence consisting of $n$ words with an $m$-word aspect, we represent it as a sequence denoted as ${\bf S}=\left\{{\bf w}_{1},{\bf w}_{2},\cdot\cdot\cdot,{\bf w}_{\gamma},\cdot\cdot\cdot,{\bf w}_{\gamma+m},\cdot\cdot\cdot,{\bf w}_{n}\right\}$, where ${\bf w}_{t}\in{\bf S}$ and $\gamma$ indicate a word token and starting index of an aspect, respectively. Each word token is then transformed into a low-dimensional vector denoted as ${\bf e}_{t}\in\mathbb{R}^{d_{w}}$, where $d_{w}$ denotes the dimension of the word embedding.
@@ 3.2 Bi-LSTM Network
Word embedding embeds the words in the sentence into low-dimensional vectors, and these embedding vectors are used as inputs to the Bi-LSTM network [11]. The utilization of Bi-LSTM allows for the extraction of contextual information from sentences. As mentioned in [12], the Bi-LSTM network comprises two components: the forward LSTM and the backward LSTM. In the forward LSTM, the hidden vector $\overrightarrow{\bf h}^{LSTM}_{t}$ at the current time step is computed by combining the input embedding ${\bf e}_{t}$ with $\overrightarrow{\bf h}^{LSTM}_{t-1}$ from the previous time step. The backward LSTM is constructed by utilizing ${\bf e}_{t}$ and $\overleftarrow{\bf h}^{LSTM}_{t+1}$ from the subsequent time step. Finally, the hidden vectors $\overrightarrow{\bf h}^{LSTM}_{t}$ and $\overleftarrow{\bf h}^{LSTM}_{t}$ are concatenated to create the overall hidden vector ${\bf h}^{LSTM}_{t}$.